OpenStudy (anonymous):

The cost of renting a car is 1500.00 per day and 50.00 per kilometer driven.Write an equation for the cost C of renting car for one day as a function of K the number of kilometers driven .If Mr.bug rented the car and drove 20 km, how much would he pay ?

OpenStudy (anonymous):

C=1500+(50*K) is the equation. Mr.Bug drove 20 Km... So, C=1500+(50*20)=2500

OpenStudy (anonymous):

thanks. :)